# Ultrafast dynamics of the Rydberg states of CO2: Autoionization and Dissociation lifetimes

### v1

@@ I Introduction
Ultrafast femtosecond and attosecond light pulses have enabled time-resolved probing of molecular dynamics. In particular, time-resolved photoelectron spectroscopy has served as an important spectroscopic tool to study photoionization and excited state dynamics in complex molecular systems [1, 2, 3]. Specifically, pump-probe time-resolved photoelectron spectroscopy has been used to gain new insights about autoionization and predissociation in Rydberg state dynamics [4, 5], non-adiabatic dynamics [6, 7, 8] and vibronic wavepacket dynamics [9, 10] in molecules.
@@
Here we report an experimental study of the Rydberg state dynamics of carbon dioxide ($CO_{2}$) molecule. It is a linear triatomic molecule with several Rydberg series of interest. Ground state electronic configuration of $CO_{2}$ is given by $[core](4\sigma_{g})^{2}(3\sigma_{u})^{2}(1\pi_{u})^{4}(1\pi_{g})^{4}$ with symmetry $X^{1}\Sigma_{g}^{+}$ [11]. The lowest three ionic continua $(1\pi_{g})^{-1}X^{2}\Pi_{g}$, $(1\pi_{u})^{-1}A^{2}\Pi_{u}$ and $(3\sigma_{u})^{-1}B^{2}\Sigma_{u}^{+}$ are at 13.778, 17.314 and 18.077 eV adiabatic ionization energies respectively [11, 12, 13, 14, 15]. The Rydberg series associated with the $A^{2}\Pi_{u}$ continuum is the Tanaka-Ogawa series [16]. The exact assignments of these series is still not very well known. The most acceptable assignment of these states are given as $(1\pi_{u})^{-1}nd\delta_{g}$. Tanaka-Ogawa Rydberg series also has seven known vibrational progressions. All these vibrational levels converge to corresponding vibrational levels of $A^{2}\Pi_{u}$ continuum [16, 17, 18]. The $B^{2}\Sigma_{u}^{+}$ continuum has two Rydberg series associated with it, namely Henning sharp $(3\sigma_{u})^{-1}nd\sigma_{g}$ and Henning diffuse $(3\sigma_{u})^{-1}ns\sigma_{g}$ series [19, 20, 16].
@@
A number of photoabsorption, photoionization and photoelectron studies have been used to study the dynamics of these Rydberg states. Between 17-18 eV excitation energy the photoexcitation spectrum is dominated by Henning sharp and diffuse series [21, 11, 22, 23]. High resolution synchrotron radiation studies have precisely measured the photoexcitation and photoionization cross-sections of these states. Recent ultrafast time resolved studies have also measured the asymmetry parameters in photoelectron distribution with branching ratios between different vibrational states [24]. These Rydberg states can undergo spontaneous preferential autoionization to $X^{2}\Pi_{g}$ continuum or they can predissociate to neutral fragments. Autoionization is an electronic correlation driven process whereas predissociation dynamics involves both nuclear and electronic motions. It is well known from theoretical quantum efficiency calculations as well as fluorescence emission spectroscopy that once these Rydberg states are excited there is a competition between autoionization and neutral predissociation [25]. Despite the extensive interest, there is very little quantitaive information available on the dissociation and autoionization lifetimes of these Rydberg states. In a recent study by Fidler et al [26] autoionization lifetimes of a few Henning diffuse [$(3\Sigma_{u})^{-1}ns\sigma_{g}$] states have been measured for the first time using four-wave-mixing spectroscopy and compared with theoretical calculations from Fano lineshapes. Another recent theoretical study by Pranjal et al [27] calculates the valence shell resonant photoionization and molecular frame photoelectron angular distributions for these states. In our experiment, we use photoelectron spectroscopy to report previously unknown predissociation and autoionization lifetimes for three Henning sharp states [$(3\sigma_{u})^{-1}nd\sigma_{g}$, n = 4, 5, 6] to give a first experimental measurement of the relaxation dynamics of these states.
@@ II Experiment
We use a pump-probe experimental setup (Fig. 1) for our photoelectron spectroscopy. We start with a titanium-sapphire laser amplifier system to generate a $\sim 50$ fs near infrared (NIR) Gaussian pulse with a central wavelength of 780 nm, pulse energy of $\sim 1.5$ mJ and repetition rate of 1 kHz. This pulse is split into two equal parts using a 50-50 beamsplitter. One half of the beam is focused into a semi-infinite gas cell filled with xenon to drive high harmonic generation (HHG) and produce attosecond extreme ultraviolet (XUV) pulse trains. This XUV pulse train is composed of only odd harmonics of the driving NIR frequency and used as the pump in our experiment. Using a toroidal mirror, XUV beam is focused onto an effusive gas jet of carbon dioxide in a velocity map imaging (VMI) spectrometer. The other half of the laser pulse is delayed with respect to the pump and focused using a 50 cm lens to the same interaction point in the VMI spectrometer. The intensity of the probe pulse at the interaction region is $\sim 1$ TW/cm2. A flat mirror with a hole in the center is used to mix the pump and the probe collinearly. Both XUV and NIR pulses are linearly polarized parallel to each other.
@@
The VMI approach [28] is a well known and effective approach in the field of photoelectron spectroscopy. In this technique, the three dimensional momentum distribution of the photoelectrons emanating from an extended focal volume is mapped onto a two dimensional position sensitive detector, thus yielding detailed information about both energy and angular distribution of the photoelectrons. The VMI spectrometer has an integrated effusive gas jet of 70 $\mu$m diameter at the center of the repeller electrode. The photoelectron momentum images are recorded using a CCD camera that images the phosphor screen placed behind the microchannel plates (MCPs). The photoelectron images at each pump-probe time delay are Abel inverted using polar basis function expansion (pBASEX) algorithm [29] to obtain 3D photoelectron momentum distribution. Angular integration of momentum distribution can be used to plot the kinetic energy distribution at each pump-probe delay to obtain time resolved spectrogram. The raw VMI images have some asymmetry due to the non-uniform spatial response of our detector. To address this, we center and symmetrize each VMI image before reconstruction. VMI spectrometer is calibrated using photoelectrons generated from the XUV photoionization of argon. For low energy electrons, small energy shifts were corrected using the photoionization lines from the Rydberg states of Xenon. The XUV spectral content is separately monitored by using a spectrometer consisting of a grating and an x-ray CCD.
@@
We subtract XUV+NIR spectrogram from XUV-only spectrogram at each time delay to extract the delay dependent effect of the probe. In the results discussed below we use this differential photoelectron spectrograms to isolate probe-induced photoionization arising from the XUV excited Rydberg states of $CO_{2}$, while also analyzing the probe induced disruption of autoionization signals.
@@
The schematic of the XUV excitation and NIR probing in our experiment is shown in the Fig. 2. The 11th harmonic in the XUV spectrum resonantly excites neutral carbon dioxide molecule from the ground state ($X^{1}\Sigma_{g}^{+}$) to a number of neutral Rydberg states between 17-18 eV. The spectrum of the 11th harmonic is peaked around 17.5 eV with a FWHM bandwidth of $\sim$0.4 eV. As mentioned before, there are three different Rydberg progressions of $CO_{2}$ that has been reported at this energy range: Henning sharp series [$(3\sigma_{u})^{-1}nd\sigma_{g}$], Henning diffuse series [$(3\sigma_{u})^{-1}ns\sigma_{g}$] and, higher vibrational states of Tanaka-Ogawa series [$(1\pi_{u})^{-1}nd\delta_{g}$)]. Both Henning sharp ($nd\sigma_{g}$) and diffuse ($ns\sigma_{g}$) series converge to $B^{2}\Sigma_{u}^{+}$ ionic continuum at 18.08 eV. Only these series are shown in Fig. 2 because they have significantly larger photoabsorption cross-section than the higher vibrational states of Tanaka-Ogawa progression. Once excited, the Rydberg states can undergo spontaneous autoionization to $X^{2}\Pi_{g}$ ionic continuum at 13.78 eV, or they can predissociate to neutral fragments $CO(A^{1}\Pi)+O(^{1}S)$. In general, autoionization and predissociation decay pathways compete, and the associated timescales ($\tau_{A}$ and $\tau_{D}$ respectively) have not been fully explored for these states. Measurement of photoionization quantum efficiency indicates that predissociation is dominant in sharp series only [16, 11]. We use a probe pulse with photon energy of $\sim 1.59$ eV at time delay $\tau$, to photoionize both the molecular Rydberg states as well as the neutral dissociation fragments available at that time delay. Thus probe induced photoionization restricts the possibility of spontaneous autoionization and we isolate the effect of the probe pulse using differential photoelectron signal as outlined above. We observe that photoionization and autoionization photoelectron angular distributions (PADs) are mutually perpendicular. This arises from the preferential autoionization direction involving change in core ion symmetry, and it allows us to identify the symmetry of participating Rydberg states. We model the photoelectron yield due to the competition between autoionization and predissociation to estimate lifetimes ($\tau_{A}$ and $\tau_{D}$) associated with two decay pathways. Finally, by fitting the time-resolved differential spectrograms we obtain the previously unknown autoionization and predissociation rates for several members of the Henning sharp Rydberg series of $CO_{2}$.
@@ III.1 Differential Spectrogram
Figure 3: Experimentally observed angle integrated time resolved differential photoelectron spectrum of the Rydberg states of $CO_{2}$. Red and blue implies the NIR probe induced enhancement or depletion of photoelectron yield. Near time delay zero we see direct electrons from XUV-NIR cross-correlation. At positive time delay we see long lived Rydberg states. Expected photoelectron kinetic energies for both autoionization and single photon photoionization are shown for Henning sharp ($nd\sigma_{g}$) and Henning difuse ($ns\sigma_{g}$) states.
@@
In Fig. 3, the time resolved differential photoelectron spectrogram is shown. At each time delay, momentum images are recorded, reconstructed and angularly integrated, to obtain kinetic energy as a function of time delay. To isolate the effect of the probe we record both (XUV+NIR) and XUV only spectra at each time delay and subtract them to get a differential spectrum. The spectrogram in Fig. 3 shows the time evolution of continuum photoelectron kinetic energies at pump-probe time delay interval of 10 fs over a delay range of -150 fs – 950 fs. Red and blue color scales correspond to positive and negative differences between (XUV+NIR) and XUV-only spectrograms, respectively. Thus, red is the enhancement and blue is the depletion in the photoelectron yield due to the effect of time delayed probe pulse. At negative time delay the NIR probe precedes the XUV pump excitation, showing no dynamics. Around time delay zero, we see single NIR-photon scattering sidebands of several direct photoionization bands of the neutral $CO_{2}$ which correspond to emergence of electrons in different ionization continua under the action of the XUV pulse train. These photoelectron bands around time delay zero correspond to XUV-NIR cross-correlation and gives information about the temporal profile of the NIR pulse.
@@
In the positive time delay side of the spectrogram we see long lived red and blue signals that correspond to the Rydberg state dynamics of the system. Once the 11th harmonic XUV pump excites neutral $CO_{2}$ to Henning sharp and diffuse series, these are further ionized to $B^{2}\Sigma_{u}^{+}$ by the delayed NIR probe pulse. The long-lived red signals between 0.6-1.5 eV kinetic energies correspond to photoelectron yield due to NIR photoionization. Correspondingly, the long-lived blue negative signals between 3-4 eV kinetic energies corresponds to the depletion of autoionization decay channel by the action of the same NIR pulse. The differential signal analysis allows us to clearly isolate the time-delayed Rydberg state dynamics by subtracting out the contribution of direct XUV photoionization. Expected kinetic energies for both photoionization and autoionization signals for different Rydberg states are shown with white tick marks and energy ladders in Fig. 3 for both Henning sharp ($nd\sigma_{g}$) and diffuse ($ns\sigma_{g}$) series. Blue autoionization signal does not show any discrete energy structure due to the fact that these Rydberg states can autoionize to multiple vibrational levels of $X^{2}\Pi_{g}$ ionic continuum thus producing different kinetic energy electrons from a single Rydberg state. On the other hand, low energy red photoionization signal shows three discrete long-lived lines.
@@
To identify the Rydberg states responsible for these photoionization lines, we use photoabsorption data from the literature [23] as shown in Fig. 4(a). A typical XUV spectrum of the $11^{th}$ harmonic is also shown. We observe that Henning sharp and diffuse series dominate the spectrum beyond $17.4$ eV. Below that energy, some of the Tanaka-Ogawa states with v = 1, 2 and 3 vibrational levels are visible with cross-sections comparable to the 4d Henning sharp state. All these series converge to their corresponding ionization limits, which are also listed in Fig. 4(a). To compare this photoabsorption data with our work, we do a baseline subtraction followed by multiplication with our XUV spectral profile, to get the scaled photoexcitation spectra relevant to our XUV pump. For each resonance in this spectrum, we obtain the expected photoelectron kinetic energy by adding the energy of single NIR photon and then subtracting the energy of its corresponding ionization limit. The expected photoelectron spectrum obtained in this fashion is shown in Fig. 4(c). For comparison, our delay-dependent photoelectron spectrogram for positive time delays ($\tau>100$ fs) is shown in Fig. 4(b). Comparing the expected and observed photoelectron kinetic energies between Fig. 4(b) and (c) it is clear that the three long lived lines in the spectrogram correspond to photoionization of 4d, 5d and 6d states from the Henning sharp series. In fact, the 4d and 5d lines are clearly isolated from the others. Photoelectrons from $n=7$ Tanaka-Ogawa states with vibrational levels v = 2 and 3 could energetically overlap with the 6d line, but as seen from the Fig. 4(c), the cross-section for 6d photoexcitation is significantly higher than the Tanaka-Ogawa state. Note that we do not see any Henning diffuse state photoelectron peaks in the spectrogram beyond 100 fs time delay because of their short lifetimes (less than 100 fs) [26]. We do not analyze the signals for time delays shorted than 100 fs, as the data gets contaminated with NIR modification of the direct XUV photoelectron bands, as discussed before. The observed 4d, 5d, and 6d Henning sharp lines match the expected positions. To extract the lifetime of the Henning sharp states, we use narrow energy windows around the sharp series as shown in Fig. 4(b) using magenta lines and shaded areas bound by those.
@@
Figure 4: (a) Photoabsorption cross-sections of neutral Rydberg states of $CO_{2}$ in the energy range 17-18 eV (adapted from [23]). Henning sharp ($nd\sigma_{g}$) and Henning diffuse ($ns\sigma_{g}$) series are shown with black and green respectively, with their convergence limit ($B^{2}\Sigma_{u}^{+}$) at 18.08 eV. Three Tanaka-Ogawa progressions are also shown for vibrational levels $v=1$ (magenta), $v=2$ (yellow) and $v=3$ (orange) with corresponding limits ($A^{2}\Pi_{u}$, $v=1,2,3$). A typical spectrum for $11^{th}$ harmonic XUV pump is shown with the green shaded area in the background. (b) Differential spectrogram for low energy photoionization signals at positive time delay ($\tau>100$ fs). Expected kinetic energies for single photon ionization of Henning sharp and diffuse states are superposed on the date. Energy windows used for extraction of lifetime and angular distribution for these states are shown with magenta lines and the bound area in between (shown with shaded rectangles). (c) Photoexcitation yield is scaled according to our broadband XUV spectrum and the Rydberg states from panel (a) are then translated to the expected continuum photoelectron energies. Photoelectron energies are calculated separately for each series by considering single NIR photon ionization of the states to the corresponding ionization limits shown in (a).

### v2

@@ I Introduction
Ultrafast femtosecond and attosecond light pulses have enabled time-resolved probing of molecular dynamics. In particular, time-resolved photoelectron spectroscopy has served as an important spectroscopic tool to study photoionization and excited state dynamics in complex molecular systems [1, 2, 3]. Specifically, pump-probe time-resolved photoelectron spectroscopy has been used to gain new insights about autoionization and predissociation in Rydberg state dynamics [4, 5], non-adiabatic dynamics [6, 7, 8], and vibronic wavepacket dynamics [9, 10] in molecules.
@@
Here we report an experimental study of the Rydberg state dynamics of carbon dioxide (CO2) molecule. It is a linear triatomic molecule with several Rydberg series of interest. Ground state electronic configuration of CO2 is given by $[core](4\sigma_{g})^{2}(3\sigma_{u})^{2}(1\pi_{u})^{4}(1\pi_{g})^{4}$ with symmetry $X^{1}\Sigma_{g}^{+}$ [11]. The lowest three ionic continua $(1\pi_{g})^{-1}X^{2}\Pi_{g}$, $(1\pi_{u})^{-1}A^{2}\Pi_{u}$ and $(3\sigma_{u})^{-1}B^{2}\Sigma_{u}^{+}$ are at 13.778, 17.314 and 18.077 eV adiabatic ionization energies, respectively [11, 12, 13, 14, 15]. The Rydberg series associated with the $A^{2}\Pi_{u}$ continuum is the Tanaka-Ogawa series [16]. The exact assignments of these series is still not very well-known. The most acceptable assignments of these states are given as $(1\pi_{u})^{-1}nd\delta_{g}$ [16]. Tanaka-Ogawa Rydberg series also has seven known vibrational progressions. All these vibrational levels converge to corresponding vibrational levels of $A^{2}\Pi_{u}$ continuum [16, 17, 18]. The $B^{2}\Sigma_{u}^{+}$ continuum has two Rydberg series associated with it, namely Henning sharp $(3\sigma_{u})^{-1}nd\sigma_{g}$ and Henning diffuse $(3\sigma_{u})^{-1}ns\sigma_{g}$ series [19, 20, 16].
@@
A number of photoabsorption, photoionization, and photoelectron studies have been used to study the dynamics of these Rydberg states. Between 17-18 eV excitation energy, the photoexcitation spectrum is dominated by Henning sharp and diffuse series. High resolution synchrotron radiation studies have precisely measured the photoexcitation and photoionization cross-sections of these states [21, 11, 22, 23]. Recent ultrafast time-resolved studies have also measured the asymmetry parameters in photoelectron distribution with branching ratios between different vibrational states [24]. These Rydberg states can undergo spontaneous preferential autoionization to the $X^{2}\Pi_{g}$ continuum or they can predissociate to neutral fragments. Autoionization is an electronic correlation driven process whereas predissociation dynamics involves both nuclear and electronic motions. It is well-known from theoretical quantum efficiency calculations as well as fluorescence emission spectroscopy that once these Rydberg states are excited there is a competition between autoionization and neutral predissociation [25]. Despite the extensive interest, there is very little quantitative information available on the dissociation and autoionization lifetimes of these Rydberg states. In a recent study by Fidler et al [26] autoionization lifetimes of a few Henning diffuse [$(3\sigma_{u})^{-1}ns\sigma_{g}$] states have been measured for the first time using four-wave-mixing spectroscopy and compared with theoretical calculations from Fano lineshapes. Another recent theoretical study by Pranjal et al [27] calculates the valence shell resonant photoionization and molecular frame photoelectron angular distributions for these states. In our experiment, we use photoelectron spectroscopy to report previously unknown predissociation and autoionization lifetimes for three Henning sharp states [$(3\sigma_{u})^{-1}nd\sigma_{g}$, n = 4, 5, 6] to give the first experimental measurement of the relaxation dynamics of these states.
@@ II Experiment
We use a pump-probe experimental setup (Fig. 1) for our photoelectron spectroscopy. We start with a titanium-sapphire laser amplifier system to generate a $\sim 50$ fs near infrared (NIR) Gaussian pulse with a central wavelength of 780 nm, pulse energy of $\sim 1.5$ mJ and repetition rate of 1 kHz. This pulse is split into two equal parts using a 50-50 beamsplitter. One half of the beam is focused into a semi-infinite gas cell filled with xenon to drive high harmonic generation (HHG) and produce attosecond extreme ultraviolet (XUV) pulse trains. This XUV pulse train is composed of only odd harmonics of the driving NIR frequency and used as the pump in our experiment. The XUV spectral content is separately monitored by using a spectrometer consisting of a grating and an x-ray CCD. Using a toroidal mirror, XUV beam is focused onto an effusive gas jet of carbon dioxide in a velocity map imaging (VMI) spectrometer. The other half of the laser pulse is delayed with respect to the pump and focused using a 50 cm lens to the same interaction point in the VMI spectrometer. The intensity of the probe pulse at the interaction region is $\sim 1$ TW/cm2. A flat mirror with a hole in the center is used to mix the pump and the probe collinearly. Both XUV and NIR pulses are linearly polarized parallel to each other.
@@
The VMI approach [28] is a well-known and effective approach in the field of photoelectron spectroscopy. In this technique, the three-dimensional momentum distribution of the photoelectrons emanating from an extended focal volume is mapped onto a two-dimensional position-sensitive detector, thus yielding detailed information about both energy and angular distribution of the photoelectrons. The VMI spectrometer has an integrated effusive gas jet of 70 $\mu$m diameter at the center of the repeller electrode. The photoelectron momentum images are recorded using a CCD camera that images the phosphor screen placed behind the microchannel plates (MCPs). The photoelectron images at each pump-probe time delay are Abel inverted using polar basis function expansion (pBASEX) algorithm [29] to obtain 3D photoelectron momentum distribution. Angular integration of momentum distribution can be used to plot the kinetic energy distribution at each pump-probe delay to obtain time resolved spectrogram. The raw VMI images have some asymmetry due to the non-uniform spatial response of our detector. To address this, we center and symmetrize each VMI image before reconstruction. VMI spectrometer is calibrated using photoelectrons generated from the XUV photoionization of argon. For low energy electrons, small energy shifts were corrected using the photoionization lines from the Rydberg states of Xenon.
@@
We subtract XUV+NIR spectrogram from XUV-only spectrogram at each time delay to extract the delay dependent effect of the probe. In the results discussed below we use this differential photoelectron spectrograms to isolate probe-induced photoionization arising from the XUV excited Rydberg states of CO2, while also analyzing the probe induced disruption of autoionization signals.
@@
The schematic of the XUV excitation and NIR probing in our experiment is shown in the Fig. 2. The 11th harmonic in the XUV spectrum resonantly excites neutral carbon dioxide molecule from the ground state ($X^{1}\Sigma_{g}^{+}$) to a number of neutral Rydberg states between 17-18 eV. The spectrum of the 11th harmonic is peaked around 17.5 eV with a FWHM bandwidth of $\sim$0.4 eV. As mentioned before, there are three different Rydberg progressions of CO2 that has been reported at this energy range: Henning sharp series [$(3\sigma_{u})^{-1}nd\sigma_{g}$], Henning diffuse series [$(3\sigma_{u})^{-1}ns\sigma_{g}$] and, higher vibrational states of Tanaka-Ogawa series [$(1\pi_{u})^{-1}nd\delta_{g}$)]. Both Henning sharp ($nd\sigma_{g}$) and diffuse ($ns\sigma_{g}$) series converge to $B^{2}\Sigma_{u}^{+}$ ionic continuum at 18.08 eV. Only these series are shown in Fig. 2 because they have significantly larger photoabsorption cross-section than the higher vibrational states of Tanaka-Ogawa progressions. Once excited, the Rydberg states can undergo spontaneous autoionization to $X^{2}\Pi_{g}$ ionic continuum at 13.78 eV, or they can predissociate to neutral fragments $CO(A^{1}\Pi)+O(^{1}S)$. In general, autoionization and predissociation decay pathways compete, and the associated timescales ($\tau_{A}$ and $\tau_{D}$ respectively) have not been fully explored for these states. Measurement of photoionization quantum efficiency indicates that predissociation is dominant in sharp series only [16, 11]. We use a probe pulse with photon energy of $\sim 1.59$ eV at time delay $\tau$, to photoionize both the molecular Rydberg states as well as the neutral dissociation fragments available at that time delay. Thus, probe induced photoionization restricts the possibility of spontaneous autoionization and we isolate the effect of the probe pulse using differential photoelectron signal as outlined above. We observe that photoionization and autoionization photoelectron angular distributions (PADs) are mutually perpendicular. This arises from the preferential autoionization direction involving change in core ion symmetry, and it allows us to identify the symmetry of participating Rydberg states. We model the photoelectron yield due to the competition between autoionization and predissociation to estimate lifetimes ($\tau_{A}$ and $\tau_{D}$) associated with two decay pathways. Finally, by fitting the time-resolved differential spectrograms we obtain the previously-unknown autoionization and predissociation rates for several members of the Henning sharp Rydberg series of CO2.
@@ III.1 Differential Spectrogram
Figure 3: Experimentally observed angle integrated time resolved differential photoelectron spectrum of the Rydberg states of CO2. Red and blue implies the NIR probe induced enhancement or depletion of photoelectron yield. Near time delay zero we see direct electrons from XUV-NIR cross-correlation. At positive time delay we see long lived Rydberg states. Expected photoelectron kinetic energies for both autoionization and single photon photoionization are shown for Henning sharp ($nd\sigma_{g}$) and Henning difuse ($ns\sigma_{g}$) states.
@@
In Fig. 3, the time resolved differential photoelectron spectrogram is shown. At each time delay, momentum images are recorded, reconstructed, and angularly integrated, to obtain kinetic energy as a function of time delay. To isolate the effect of the probe we record both (XUV+NIR) and XUV only spectra at each time delay and subtract them to get a differential spectrum. The spectrogram in Fig. 3 shows the time evolution of continuum photoelectron kinetic energies at pump-probe time delay interval of 10 fs over a delay range of -150 fs – 950 fs. Red and blue color scales correspond to positive and negative differences between (XUV+NIR) and XUV-only spectrograms, respectively. Thus, red is the enhancement and blue is the depletion in the photoelectron yield due to the effect of time delayed probe pulse. At negative time delay, the NIR probe precedes the XUV pump excitation, showing no dynamics. Around time delay zero, we see single NIR-photon scattering sidebands of several direct photoionization bands of the neutral CO2 which correspond to emergence of electrons in different ionization continua under the action of the XUV pulse train. These photoelectron bands around time delay zero correspond to XUV-NIR cross-correlation and gives information about the temporal profile of the NIR pulse.
@@
In the positive time delay side of the spectrogram, we see long lived red and blue signals that correspond to the Rydberg state dynamics of the system. Once the 11th harmonic XUV pump excites neutral CO2 to Henning sharp and diffuse series, these are further ionized to $B^{2}\Sigma_{u}^{+}$ by the delayed NIR probe pulse. The long-lived red signals between 0.6-1.5 eV kinetic energies correspond to photoelectron yield due to NIR photoionization. Correspondingly, the long-lived blue negative signals between 3-4 eV kinetic energies represent the depletion of autoionization decay channel by the action of the same NIR pulse. The differential signal analysis allows us to clearly isolate the time-delayed Rydberg state dynamics by subtracting out the contribution of direct XUV photoionization. Expected kinetic energies for both photoionization and autoionization signals for different Rydberg states are shown with white tick marks and energy ladders in Fig. 3 for both Henning sharp ($nd\sigma_{g}$) and diffuse ($ns\sigma_{g}$) series. Blue autoionization signal does not show any discrete energy structure due to the fact that these Rydberg states can autoionize to multiple vibrational levels of $X^{2}\Pi_{g}$ ionic continuum thus producing different kinetic energy electrons from a single Rydberg state. On the other hand, low energy red photoionization signal shows three discrete long-lived lines.
@@
To identify the Rydberg states responsible for these photoionization lines, we use photoabsorption data from the literature [23] as shown in Fig. 4(a). A typical XUV spectrum of the $11^{th}$ harmonic is also shown. We observe that Henning sharp and diffuse series dominate the spectrum beyond $17.4$ eV. Below that energy, some of the Tanaka-Ogawa states with v = 1, 2 and 3 vibrational levels are visible with cross-sections comparable to the 4d Henning sharp state. All these series converge to their corresponding ionization limits, which are also listed in Fig. 4(a). To compare this photoabsorption data with our work, we do a baseline subtraction followed by multiplication with our XUV spectral profile, to get the scaled photoexcitation spectra relevant to our XUV pump. For each resonance in this spectrum, we obtain the expected photoelectron kinetic energy by adding the energy of single NIR photon and then subtracting the energy of its corresponding ionization limit. The expected photoelectron spectrum obtained in this fashion is shown in Fig. 4(c). For comparison, our delay-dependent photoelectron spectrogram for positive time delays ($\tau>100$ fs) is shown in Fig. 4(b). Comparing the expected and observed photoelectron kinetic energies between Fig. 4(b) and (c) it is clear that the three long-lived lines in the spectrogram correspond to photoionization of 4d, 5d, and 6d states from the Henning sharp series. In fact, the 4d and 5d lines are clearly isolated from the others. Photoelectrons from $n=7$ Tanaka-Ogawa states with vibrational levels v = 2 and 3 could energetically overlap with the 6d line, but as seen from the Fig. 4(c), the cross-section for 6d photoexcitation is significantly higher than the Tanaka-Ogawa state. Note that we do not see any Henning diffuse state photoelectron peaks in the spectrogram beyond 100 fs time delay because of their short lifetimes (less than 100 fs) [26]. We do not analyze the signals for time delays shorted than 100 fs, as the data gets contaminated with NIR modification of the direct XUV photoelectron bands, as discussed before. The observed 4d, 5d, and 6d Henning sharp lines match the expected positions. To extract the lifetime of the Henning sharp states, we use narrow energy windows around the sharp series as shown in Fig. 4(b) using magenta lines and shaded areas bound by those.
@@
Figure 4: (a) Photoabsorption cross-sections of neutral Rydberg states of CO2 in the energy range 17-18 eV (adapted from [23]). Henning sharp ($nd\sigma_{g}$) and Henning diffuse ($ns\sigma_{g}$) series are shown with black and green respectively, with their convergence limit ($B^{2}\Sigma_{u}^{+}$) at 18.08 eV. Three Tanaka-Ogawa progressions are also shown for vibrational levels $v=1$ (magenta), $v=2$ (yellow) and $v=3$ (orange) with corresponding limits ($A^{2}\Pi_{u}$, $v=1,2,3$). A typical spectrum for $11^{th}$ harmonic XUV pump is shown with the green shaded area in the background. (b) Differential spectrogram for low energy photoionization signals at positive time delay ($\tau>100$ fs). Expected kinetic energies for single photon ionization of Henning sharp and diffuse states are superposed on the date. Energy windows used for extraction of lifetime and angular distribution for these states are shown with magenta lines and the bound area in between (shown with shaded rectangles). (c) Photoexcitation yield is scaled according to our broadband XUV spectrum and the Rydberg states from panel (a) are then translated to the expected continuum photoelectron energies. Photoelectron energies are calculated separately for each series by considering single NIR photon ionization of the states to the corresponding ionization limits shown in (a).